OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(z - 2) = 125?

OpenStudy (anonymous):

take logs of both sides: ln 25^(z - 2) = ln 125 (z - 2) ln 25 = ln 125 - using laws of logarithms z-2 = ln 125/ ln25 can you finish this off?

OpenStudy (anonymous):

yeah I think I got it z=3.5 right?

OpenStudy (anonymous):

checking... yea

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ty